# Seeing Through Deception: Uncovering Misleading Creator Intent in Multimodal News with Vision-Language Models

### v2

@@ 3.5 Evaluation Tasks and Metrics
DeceptionDecoded facilitates the evaluation of the following three intent-centric tasks:
  • Task 1: Misleading Intent Detection. Given an input triplet of either $\{I^{\prime},T,A\}$ or $\{I,T^{\prime},A\}$, the goal is to predict whether the instance is misleading (M) or non-misleading (NM).
  • Task 2: Misleading Source Attribution. A 3-way classification task in which the model must attribute the source of misleading intent to either the image or the text modality, or output “NA” if the instance is non-misleading.
  • Task 3: Creator Desire Inference. A multi-label classification task where the model must identify the creator’s intended societal influence, selecting up to three options from a predefined list (see Appendix A.3).
We evaluate Tasks 1 and 2 using classification accuracy, and Task 3 using F1 score.
@@ 4 Experiments
Table 2: VLM accuracy (%) on misleading intent detection (black) and misleading source attribution (colored). I refers to the implication-oriented approach, and C refers to the consistency-oriented approach (see Section 4.1). Green cells indicate the top-2 best-performing models. Red cells indicate hallucinations observed in smaller VLMs, where the model attributes a misleading source (M) despite predicting the instance as non-misleading (NM).
@@
Table 3: Performance (F1%) of VLMs on creator desire inference. Green cells indicate the best-performing models, while red cells indicate models that failed to make a single correct prediction.
4.1 Experimental Details
With DeceptionDecoded, we evaluate 14 representative vision-language models (VLMs) spanning a range of sizes, model families, and access levels. These include (1) multimodal Large Reasoning Models (MLRMs), such as Claude-3.7-Sonnet Anthropic (2025c); (2) proprietary multimodal Large Language Models (MLLMs), such as GPT-4o Hurst et al. (2024); and (3) open-source MLLMs, such as the Qwen-2.5-VL series Bai et al. (2025). The complete list of models used in our experiments is provided in Table 7.
We evaluate two representative paradigms in misinformation detection, distinguished by their reasoning focus: (1) The implication-oriented (I) approach, which focuses on the inferred implications of the news sample; and (2) The consistency-oriented (C) approach, which assesses consistency both across modalities and between the image–caption pair and the reference article. Unless otherwise specified, all experiments on DeceptionDecoded default to the consistency-oriented (C) approach. Experimental setup and evaluation prompt details are shown in Appendix C.1.
4.2 Main Results
Consistency-Oriented Reasoning Generally Outperforms Implication-Oriented Reasoning. As shown in Table 2, a possible reason is that misleading news is often constructed through deliberate content manipulation. Detecting unsubstantiated inconsistencies, either between the image and its caption or between the image-caption pair and the reference article, helps models better identify misleading intent.
@@ 4.2 Main Results
Table 4: Misleading intent detection performance of VLMs when exposed to spurious authenticity-related hints, averaged over both misleading and non-misleading cases. Hint construction details are provided in Section 5.1.
Model M-Sub (Text) M-Sig (Text)
I+T T+A Full I+T T+A Full
GPT-4o 21.8 63.4 70.2 51.6 88.4 93.3
GPT-4o-mini 3.6 36.8 18.8 10.2 64.4 45.9
Qwen2.5-72B 10.6 53.6 49.6 24.6 83.8 82.5
Qwen2.5-32B 6.2 35.8 27.5 18.4 67.2 58.6
Table 5: VLM performance (Acc. %) on misleading intent prediction under partial modality settings, suggesting that smaller VLMs may overly rely on image-text consistency for misleading intent detection. (I+T: image & caption only; T+A: text & reference article only; Full: all modalities provided.)
Figure 3: Misleading intent detection performance of VLMs on misleading text presented either in a professional tone (as used in the original DeceptionDecoded setting) or in an explicitly misleading style.
VLMs Struggle to Reason About Misleading Creator Intent.
As shown in Table 2, even state-of-the-art MLRMs such as Claude-3.7 demonstrate limited performance in detecting and attributing misleading creator intent. Performance is even lower on the more complex task of creator desire inference (Table 3), suggesting a deeper challenge in reasoning about communicative goals.
These findings raise alarming concerns. Our setup reflects a realistic deployment scenario, where readers are exposed to subtle yet misleading representations of otherwise trustworthy multimodal news and must rely on automated systems for guidance. The consistent underperformance of VLMs under such conditions suggests an urgent need to investigate the underlying patterns these models exploit when assessing content authenticity, inspiring our in-depth analysis in Section 5.
5 Analysis
5.1 Probing the Limitations of VLMs
VLMs Can Be Misled by Surface-Level Image–Text Consistency.
Although misleading intent is often not substantiated when compared with the trustworthy reference article, the manipulated image and caption typically appear internally consistent, as illustrated in Figure 1. Ideally, models should integrate information across all modalities (i.e., image, text, and article) to detect such deceptive intent. To investigate this, we evaluate VLMs under two partial input settings: (1) Image + Text, and (2) Text + Article.
However, results in Table 5 reveal a contradiction to this ideal. For instance, GPT-4o-mini performs better when detecting misleading text by comparing it with the article, but its performance drops notably when the misleading image is added. This suggests that consistency between image and text can mask deception, misleading models that lack deeper reasoning about cross-source veracity.
VLMs Fail to Detect Misleading Text Framed in Credible-Sounding Styles.
Prior work in textual misinformation detection Chen and Shu (2024); Wu et al. (2024) has shown that large language models (LLMs) often rely on stylistic cues rather than factual correctness when making veracity judgments. To investigate whether this stylistic bias extends to the multimodal setting, we reframe misleading captions in DeceptionDecoded from professional-sounding styles to more overtly deceptive or sensational tones (see Appendix C.2). As illustrated in Figure 3, we observe a similar trend: VLMs are more likely to misclassify misleading content when it is framed in a polished, authoritative style. This suggests that current models may be overly influenced by surface-level linguistic cues, posing a serious risk when misinformation is intentionally crafted to appear credible.
Models React Strongly to Spurious Authenticity Cues in Prompts.
@@ 5.1 Probing the Limitations of VLMs
To examine whether VLMs are influenced by high-level framing cues that are not grounded in the input content, we inject two types of authenticity-related hints into the prompt (see Appendix C.3): (1) Skeptical, which presumes the news piece may contain intentional distortions. This hint serves as a helpful signal for misleading samples but becomes adversarial for non-misleading samples; and (2) Trusting, which assumes the content is reliable, reversing the helpful and adversarial roles.
@@
As shown in Table 4, these prompts produce significant and consistent performance changes: they improve accuracy in helpful cases and degrade performance in adversarial ones. This reveals that VLMs are susceptible to spurious prior assumptions about content authenticity, even when such cues are not supported by the input.
5.2 Impact of Image Generation Advances
DeceptionDecoded includes reasonably high-quality images generated using the FLUX model, as supported by human validation results in Section 3.4. As image generation capabilities continue to advance, we conduct a pilot study to assess whether VLM performance improves when exposed to higher-fidelity images generated by GPT-image-1 OpenAI (2025a), using a subset of 200 sampled misleading instances.
As shown in Table 6 and Figure 4, we observe modest performance gains, likely due to the more vivid visual portrayal of creator intent. However, despite these improvements, understanding the underlying communicative intent behind multimodal news remains a significant challenge even when image realism is enhanced. These findings highlight that surface-level fidelity alone is insufficient, reinforcing the need for models that reason explicitly about creator objectives.
The growing realism of deceptive images as reflected by GPT-generated instances introduces a significant risk. When high-fidelity images convincingly align with misleading textual narratives, the deception becomes more subtle and harder to detect. These advances further underscore the importance of reasoning about the creator’s underlying intent, rather than relying solely on content realism or modality alignment.
Figure 4: Case Study: a VLM (GPT-4o-mini) fails to detect misleading creator intent in a FLUX-generated image due to its inability to recognize the unsubstantiated presence of a crowd with fireworks, given the news context. The state-of-the-art GPT-image-1 model produces a clearer depiction, leading to a correct prediction.
@@ 5.2 Impact of Image Generation Advances
Table 6: Misleading intent detection performance of VLMs on images created using different image generation models, evaluated on a pilot test set of 200 samples. While performance improves with higher-fidelity GPT images, a substantial gap remains.
@@ 6 Conclusion
In this work, we establish the important yet underexplored task of understanding creator intent in multimodal misinformation. We introduce DeceptionDecoded, a large-scale benchmark designed to evaluate VLMs across intent-centric tasks. Through extensive experiments, we uncover alarming vulnerabilities in state-of-the-art VLMs, which over-rely on spurious shortcuts such as cross-modal consistency, stylistic cues, and heuristic indicators of authenticity to detect misleading intent. Our findings pave the way for future research on MMD systems capable of intent-aware, authenticity-centric semantic reasoning toward effective misinformation governance.
Limitations
@@
Through DeceptionDecoded, we establish a framework for understanding misleading creator intent in multimodal news, providing a meaningful first step toward modeling and detecting misleading creator intent. While DeceptionDecoded presents a structured and scalable framework, it also opens up several avenues for further research in advancing intent-aware misinformation governance across more realistic and socially grounded scenarios.
@@
First, although our framework supports the evaluation of creator intent interpretation through tasks such as misleading intent detection and creator desire inference, fully open-ended intent articulation and corresponding evaluation remains an open challenge. Future work could explore adaptive prompting strategies and leverage the emerging LLM-as-a-Judge paradigm Gu et al. (2024) to assess generated intent representations in more flexible and scalable ways.
@@
Second, our current benchmark focuses on intent detection and interpretation, but does not directly measure downstream societal influence. In real-world scenarios, the effectiveness of misinformation depends not only on creator intent but also on how audiences interpret and react to the content. Incorporating social grounding, such as user feedback signals, comment chains, or engagement patterns on social media, would provide richer context for evaluating the influence of misleading content and calibrating model responses accordingly.
Despite these limitations, DeceptionDecoded represents an important contribution toward systematic, intent-aware evaluation of VLMs in real-world misinformation contexts. We hope our benchmark and insights will serve as a solid foundation for future work.
Ethics Statement
@@
The intent-guided simulation of multimodal misinformation in DeceptionDecoded involves strategies that could potentially be misused to generate misleading content using VLMs and open-source image generation models (e.g., FLUX). Nonetheless, given the realistic and challenging nature of detecting misleading creator intent in multimodal news, we believe it is important to transparently acknowledge these risks while reporting empirical findings on the performance and limitations of current state-of-the-art VLMs.
@@
Our benchmark is developed solely to highlight the important yet underexplored aspect of modeling malicious creator intent in MMD, and to advance the understanding of current system limitations.To minimize harm, we explicitly avoid manipulating news content involving specific, identifiable individuals; all simulated content is based on anonymized or synthetic entities. To further mitigate misuse, we will open-source the dataset and evaluation scripts, but refrain from releasing the specific generation prompts that could be repurposed for deceptive use. Data access will be granted only to verified researchers under a binding usage agreement to ensure responsible use. All data collection and usage comply with the terms of service of the underlying models and platforms.
Appendix A DeceptionDecoded Details
A.1 News Topics
@@
VisualNews Liu et al. (2021) provides fine-grained topic annotations for each news article. For DeceptionDecoded, we select 10 major topic categories based on sample abundance, topical diversity, and real-world relevance.
@@
A.2 Source Data Acquisition

### v3

@@ 3.5 Evaluation Tasks and Metrics
We design three intent-centric tasks in DeceptionDecoded to evaluate complementary dimensions of creator intent reasoning.
  • Task 1: Misleading Intent Detection. Given an input triplet of either $\{I^{\prime},T,A\}$ or $\{I,T^{\prime},A\}$, the objective is to predict whether the instance is misleading (M) or non-misleading (NM).
  • Task 2: Misleading Source Attribution. A three-way classification task in which the model must attribute the source of misleading intent to either the image or the text modality, or output “NA” if the instance is non-misleading.
  • Task 3: Creator Desire Inference. A multi-label classification task where the model identifies the creator’s intended societal influence, selecting up to three options from a predefined list (see details in Appendix D.3).
Evaluation for Tasks 1 and 2 is based on classification accuracy, and for Task 3 on F1 score.
@@ 4 Experiments
Table 2: VLM accuracy (%) on misleading intent detection (black) and source attribution (colored). I denotes the implication-oriented approach, and C the consistency-oriented approach (see Section 4.1). Green cells mark the top-2 models. Red cells highlight hallucinations in smaller VLMs, where a misleading source (M) is attributed despite the instance being predicted as non-misleading (NM).
4.1 Experimental Setup
Using DeceptionDecoded, we evaluate 14 representative vision-language models (VLMs) spanning a range of sizes, model families, and access levels. These include (1) multimodal Large Reasoning Models (MLRMs), such as Claude-3.7-Sonnet (Anthropic, 2025b); (2) proprietary multimodal Large Language Models (MLLMs), such as GPT-4o (Hurst et al., 2024); and (3) open-source MLLMs, such as the Qwen-2.5-VL series (Bai et al., 2025). The complete list of models used in our experiments is provided in Table 8.
We evaluate two paradigms in misinformation detection, distinguished by their reasoning focus: (1) The implication-oriented (I) approach, which focuses on the inferred implications of the news sample; and (2) The consistency-oriented (C) approach, which assesses consistency both across modalities and between the image–caption pair and the trustworthy news context. Unless otherwise specified, all experiments on DeceptionDecoded default to the consistency-oriented (C) approach. Details of the experimental setup and evaluation prompts are provided in Appendix F.1.
4.2 Main Results
Consistency-Oriented Reasoning Outperforms Implication-Oriented Reasoning. As shown in Table 2, the consistency-oriented approach generally yields stronger performance than the implication-oriented approach. A likely explanation is that misleading news is often created through deliberate content manipulation. Detecting unsubstantiated inconsistencies—either between the image and its caption or between the image–caption pair and the trustworthy news context—provides models with a more reliable signal for identifying misleading intent.
VLMs Struggle to Reason About Misleading Creator Intent. As shown in Table 2, even state-of-the-art MLRMs such as Claude-3.7 exhibit limited performance in detecting and attributing misleading creator intent. Performance drops further on the more challenging task of creator desire inference (Table 4), indicating a deeper difficulty in reasoning about underlying communicative goals.
These findings raise concerning implications. Our evaluation reflects realistic deployment scenarios in which readers encounter subtle yet misleading representations of otherwise trustworthy multimodal news and must rely on automated systems for support. The consistent underperformance of VLMs under such conditions highlights the urgent need to examine the features and heuristics these models rely on when assessing authenticity, motivating our in-depth analyses presented in Section 4.3.
@@ 4.2 Main Results
Table 3: Performance (F1%) on creator desire inference. Green cells indicate the best-performing models, and red cells zero performance.
Model M-Sub (Text) M-Sig (Text)
I+T T+A Full I+T T+A Full
GPT-4o 21.8 63.4 70.2 51.6 88.4 93.3
GPT-4o-mini 3.6 36.8 18.8 10.2 64.4 45.9
Qwen2.5-72B 10.6 53.6 49.6 24.6 83.8 82.5
Qwen2.5-32B 6.2 35.8 27.5 18.4 67.2 58.6
Table 4: VLM performance (Acc.%) on misleading intent prediction under partial modality settings. Results suggest that smaller VLMs tend to over-rely on image–text consistency when detecting misleading intent. (I+T: image & caption only; T+A: caption & reference article only; Full: all modalities.)
4.3 VLMs Struggle Beyond Surface-Level Signals
VLMs Can Be Misled by Image–Text Semantic Consistency. Although misleading intent is often unsupported when compared with the trustworthy news context, the manipulated image and caption typically appear internally consistent (Figure 1). Ideally, models should integrate information across all three modalities: image, text, and article, to identify such deception.
Figure 3: Misleading intent detection performance of VLMs on misleading text presented either in a professional tone (as used in the original DeceptionDecoded setting) versus an explicitly misleading style.
To investigate this, we evaluate VLMs under two partial input settings: (1) Image + Text, and (2) Text + Article. As shown in Table 4, the stronger GPT-4o model follows the expected trend, achieving its best performance when all modalities are available. In contrast, smaller VLMs deviate from this ideal. For example, GPT-4o-mini performs better when detecting misleading text by comparing it directly with the article, yet its performance drops notably once the misleading image is added. This suggests that internal image–text consistency can mask deception, exposing a weakness in models that lack deeper reasoning about the credibility of information across modalities.
VLMs Fail to Detect Misleading Text Framed in Credible-Sounding Styles. Prior work on textual misinformation detection (Chen & Shu, 2024; Wu et al., 2024) shows that large language models (LLMs) often rely on stylistic cues rather than factual correctness when making veracity judgments. To test whether this bias extends to the multimodal setting, we reframe misleading captions in DeceptionDecoded from professional, authoritative tones to more explicitly deceptive or sensational ones (see details in Appendix F.2). As illustrated in Figure 3, we observe the same trend: VLMs are more likely to misclassify misleading content when it is presented in a polished, credible style. This suggests that current models may be disproportionately influenced by surface-level linguistic cues, posing serious risks when misinformation is deliberately crafted to appear authoritative.
@@ 4.3 VLMs Struggle Beyond Surface-Level Signals
Table 5: VLM performance on misleading intent detection with spurious authenticity hints, averaged over misleading and non-misleading cases. Helpful / Adversarial hint construction details are provided in Section 4.3.
@@
VLMs React Strongly to Spurious Authenticity Cues in Prompts. We further test whether models are influenced by high-level framing cues that are not grounded in the input content. To do so, we inject two types of authenticity-related hints into the prompts (see Appendix F.3): (1) Skeptical, which presumes the news piece may contain intentional distortions (helpful for misleading samples but adversarial for non-misleading ones), and (2) Trusting, which assumes the content is reliable (reversing the helpful/adversarial roles).
@@
As shown in Table 5, these hints produce significant and systematic performance shifts: they improve accuracy in helpful cases but sharply degrade it in adversarial ones. This demonstrates that VLMs remain highly sensitive to spurious prior assumptions about authenticity, even when such cues are not supported by the input.
5 Discussion
5.1 Broader Utility of Creator Intent Identification
Model Macro-F1 (%) Macro-Acc. (%)
Llava-v1.6-7B 44.41 50.31
   + FFT (w/ 6k samples) 52.37 (+7.96) 63.52 (+13.21)
Qwen2.5-VL-7B 27.96 52.12
   + FFT (w/ 6k samples) 58.66 (+30.70) 67.24 (+15.12)
Table 6: Further utility of DeceptionDecoded in enhancing multimodal misinformation detection (MMD). Fine-tuning VLMs on DeceptionDecoded yields consistent performance improvements on MMFakeBench, a general MMD benchmark.
A central contribution of DeceptionDecoded is its focus on misleading creator intent, an aspect largely overlooked in existing multimodal misinformation benchmarks yet essential for developing robust MMD systems. To examine its broader utility, we conduct a transfer experiment on the MMFakeBench validation set (Liu et al., 2025), which contains 1,000 samples (700 fake, 300 real). Specifically, we compare the performance of LLaVA-v1.6-7B and Qwen-2.5-VL-7B before and after full-parameter fine-tuning on 6,000 DeceptionDecoded samples (1,000 per class across six intent categories, with misleading samples grouped as “Fake” and non-misleading as “Real”) for one epoch. The setup is detailed in Appendix F.5. As shown in Table 6, fine-tuning yields consistent and significant performance gains across metrics, indicating that DeceptionDecoded not only facilitates intent-based modeling but also enhances general MMD.
5.2 Image Realism Outpaces VLMs’ MMD Capabilities
DeceptionDecoded incorporates reasonably high-quality images generated with FLUX.1 [dev] (Black Forest Labs, 2024), as confirmed by human validation in Section 3.4 and Appendix E. To examine the implications of recent advances in image generation, we conducted a controlled evaluation using a subset of 200 sampled misleading instances, replacing FLUX outputs with higher-fidelity versions generated by GPT-image-1 (OpenAI, 2025b).
@@ 5.2 Image Realism Outpaces VLMs’ MMD Capabilities
Table 7: Misleading intent detection performance of VLMs on images generated by different models. Evaluation is conducted on a 200-sample pilot set.
As shown in Table 7 and Figure 4, several VLMs exhibit modest, consistent gains when evaluated on GPT-generated images, likely because higher visual fidelity clarifies intent-related cues (for example, more legible, vivid depictions of fireworks and crowds than the FLUX-generated variant). These observations indicate that stronger visuals can make certain manipulations easier to detect. However, despite these improvements, modeling the underlying communicative intent in multimodal news remains difficult. Surface-level fidelity is insufficient; effective detection requires reasoning explicitly about creator objectives.
At the same time, the growing realism of GPT-generated images introduces substantial risks. When high-fidelity visuals align convincingly with misleading textual narratives, the resulting deception becomes more subtle and more difficult to detect. In our experiments, prompting GPT-image-1 to simulate malicious creators, without explicitly requesting misinformation, is able to bypass the established safety guardrail and generate convincing manipulations for all 200 cases. The combination of scalability and realism heightens risk, particularly since VLMs remain weak at intent detection even under these enhanced conditions.
Figure 4: Case Study: a VLM (GPT-4o-mini) fails to detect misleading creator intent in a FLUX-generated image. The model overlooks the presence of a crowd with fireworks, an unsubstantiated addition given the powerhouse fire described in both the news context and the caption. In contrast, the state-of-the-art GPT-image-1 model produces a more vivid depiction of fireworks, enabling the correct prediction.
@@ 6 Conclusion
We establish creator intent as a central perspective for multimodal misinformation and introduce DeceptionDecoded, a large-scale benchmark grounded in trustworthy news contexts. Through extensive experiments on 14 VLMs, we reveal that current state-of-the-art models are weak at intent reasoning, often relying on spurious cues such as cross-modal consistency, stylistic polish, and heuristic authenticity signals. These findings demonstrate that advances in visual fidelity are insufficient on their own, highlighting the need for authenticity-centric, intent-aware modeling. Taken together, DeceptionDecoded offers not only an evaluation benchmark but also frames a research agenda, opening new directions for intent-aware evaluation and governance of multimodal misinformation at scale.  [^1] Limitations and ethical considerations are discussed in Appendix A and Appendix B.
Appendix A Limitations and Future Work
@@
Through DeceptionDecoded, we establish a structured and scalable framework for understanding misleading creator intent in multimodal news, representing a meaningful first step toward intent-aware multimodal misinformation governance. While our benchmark makes a substantial contribution, it also opens several promising directions for extending its value and impact.
@@
First, although DeceptionDecoded supports intent interpretation through tasks such as misleading intent detection and creator desire inference, fully open-ended intent articulation and evaluation remains an open challenge. Future work could explore adaptive prompting strategies or leverage emerging paradigms such as LLM-as-a-Judge (Gu et al., 2024) to assess generated intent representations in more flexible, reliable, and scalable ways.
@@
Second, while our current focus is on intent detection and interpretation, the downstream societal influence of misinformation is not directly measured. In real-world settings, the effectiveness of misleading narratives depends not only on creator intent but also on how audiences interpret and respond. Incorporating social grounding (such as user feedback signals, comment chains, or engagement patterns) would provide richer context for evaluating the reach and impact of misinformation, and for calibrating model responses accordingly.
Overall, DeceptionDecoded provides a principled foundation for systematic, intent-aware evaluation of VLMs, while also charting a path toward more socially grounded, resilient, and scalable approaches for governing multimodal misinformation.
Appendix B Ethics Statement
@@
The intent-guided simulation of multimodal misinformation in DeceptionDecoded necessarily involves strategies that could be misused to generate misleading content with VLMs and open-source image generation models (e.g., FLUX). Nonetheless, given the pressing challenge of detecting misleading creator intent in multimodal news, it is important to transparently acknowledge these risks while reporting empirical findings on the strengths and limitations of current state-of-the-art VLMs.
@@
Our benchmark is developed solely to highlight the important yet underexplored aspect of modeling malicious creator intent in MMD, and to advance the understanding of current system limitations. To minimize harm, we explicitly avoid manipulating news content involving specific, identifiable individuals; all simulated content is based on anonymized or synthetic entities. To further mitigate misuse, we will open-source the dataset and evaluation scripts but refrain from releasing the specific generation prompts that could be repurposed for deception. Access will be restricted to verified researchers under a binding usage agreement, and all data collection complies with the terms of service of the underlying models and platforms. Overall, we prioritize transparency while enforcing safeguards against misuse.
Importantly, our work does not seek to promote or enforce a single interpretation of news events. Instead, it focuses on detecting creator intent when multimodal content distorts the implications of a trustworthy news context by embedding unsubstantiated claims or implications (see example in Figure 1 and Figure 2). This framing distinguishes between responsible reporting, which preserves neutrality, and misleading reporting, which introduces unsupported interpretations. Ultimately, our goal is to support the development of intent-aware systems that safeguard trustworthy reporting while respecting the rights of information consumers to hold diverse viewpoints.
Appendix C LLM Usage Statement
LLMs were used solely for minor grammatical editing of the manuscript. All methodological design, analysis, and writing were conducted under full human oversight and responsibility.
Appendix D DeceptionDecoded Details
D.1 News Topics
@@
VisualNews (Liu et al., 2021) provides fine-grained topic annotations for each news article. For DeceptionDecoded, we select 10 major topic categories based on sample abundance, topical diversity, and real-world relevance.
@@
D.2 Source Data Acquisition